# FMPlug: Plug-In Foundation Flow-Matching Priors for Inverse Problems

### v1

@@ 1 Introduction
Traditional ideas for solving IPs rely on optimization formulations and associated numerical optimization methods, often motivated under the Maximum A Posteriori (MAP) estimation principle:
@@
Here, minimizing data fitting loss $\ell(\bm{y},\mathcal{A}(\bm{x}))$ encourages $\bm{y}\approx\mathcal{A}(\bm{x})$, and the regularization term $\Omega(\bm{x})$ encodes prior knowledge of ideal solutions to resolve ambiguities and hence mitigate potential ill-posedness. Advances in deep learning (DL) have revolutionized IP solving. Different DL-based approaches for IPs operate with different data-knowledge tradeoffs. For example, supervised approaches take paired datasets $\left\{\bm{y}_{i},\bm{x}_{i}\right\}_{i=1}^{N}$ and directly learn the inverse mapping, with or without using $\mathcal{A}$ [1, 2, 3]; alternative, data-driven priors learned from object-only datasets $\left\{\bm{x}_{i}\right\}_{i=1}^{N}$ can be integrated with Eq. 1.1 [4, 5, 6]; strikingly, untrained DL models themselves can serve as plug-in priors for Eq. 1.1, without any extra data [7, 8, 9, 10, 11, 12]. [1] give comprehensive reviews of these ideas.
In this paper, we focus on solving IPs with pretrained deep generative priors (DGPs) [4]. Compared to supervised approaches that perform task-specific training, these approaches enjoy great flexibility, as DGPs can be plugged in and reused for different tasks related to the same family of objects. Among the different DGPs, we are most interested in those based on diffusion models (DM), and the more general flow-matching (FM) models [13], as they are the backbone of state-of-the-art (SOTA) generative models in several domains [14, 15, 16].
Numerous recent works target solving IPs with pretrained DM/FM priors [5]. Although promising, most of them are based on domain-specific DM/FM priors, e.g., trained on the FFHQ dataset for human faces and the LSUN bedrooms dataset for bedroom scenes. This limits the practicality of these methods, as domain-specific DM/FM models are not always readily available. On the other hand, the emergence of domain-agnostic foundation DM/FM models, such as Stable Diffusion [15] and Flux [14] for images, seems to obsolete domain-specific developments; [17, 18, 19] propose such ideas. However, the reported performance clearly lags behind those obtained with domain-specific FM/DM priors, and even behind those obtained with untrained priors; see Section 2.3. This should not be surprising, as in terms of solving any particular IP, foundation priors are considerably weaker than domain-specific ones.
@@ 2.1 Flow matching (FM)
Flow Matching (FM) models are an emerging class of deep generative models [13]. They learn a continuous flow to transform a prior distribution $p_{0}(\bm{z})$ into a target distribution $p_{1}(\bm{x})$—in the same spirit of continuous normalizing flow (CNF), described by an ordinary differential equation (ODE)
@@
DMs based on probability flow ODEs can also be interpreted as FMs, although they match the score functions induced by the chosen probability path, not the vector field [13]. So, FM can be viewed as a general generative framework that covers DM as well besides other algorithmic possibilities.
@@ 2.2 Pretrained FM priors for IPs
Recent methods that use pretrained FM priors for solving IPs can be classified into two families: (1) The interleaving approach interleaves the generation steps (i.e., numerical integration steps) with gradient steps toward feasibility (i.e., making $\bm{x}$ satisfies that $\bm{y}\approx\mathcal{A}(\bm{x})$[20, 17, 18]; see a sample algorithm in Algorithm 1. Despite the simplicity and empirical effectiveness on simple IPs [20, 17, 18], these methods might not converge or return an $\bm{x}$ that respects the data prior (i.e., manifold feasibility) and/or satisfies the measurement constraint $\bm{y}\approx\mathcal{A}(\bm{x})$ (i.e., measurement feasibility); and (2) the plug-in approach views the entire generation process (lines 4–8 in Algorithm 2) as a function $\mathcal{G}_{\bm{\theta}}$ that maps any source sample to a target sample, and plugs the prior into Eq. 1.1 to obtain a unified optimization formulation [19, 6, 21] ($\circ$: function composition):
$\bm{z}^{\ast}\in\operatorname*{arg\,min}\nolimits_{\bm{z}}\;\mathcal{L}(\bm{z})\doteq\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}+\Omega\circ\mathcal{G}_{\bm{\theta}}(\bm{z});$ (2.3)
see Algorithm 2 for a sample algorithm. The estimated object is then $\mathcal{G}_{\bm{\theta}}(\bm{z}^{\ast})$. Here, the generator $\mathcal{G}_{\bm{\theta}}$ is fixed and the output $\mathcal{G}_{\bm{\theta}}(\bm{z})$ naturally satisfies the manifold feasibility. In addition, global optimization of $\mathcal{L}(\bm{z})$ forces small ($)\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}$, and hence $\bm{y}\approx\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})$, i.e., measurement feasibility.
@@
5:  $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$triangleright learned velocity
6:  $\bm{z}^{\prime}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$triangleright discrete integration
7:  $\bm{z}_{i+1}\leftarrow$ $(\bm{y},\mathcal{A})$-driven update of $\bm{z}^{\prime}_{i+1}$triangleright reducing $\ell(\bm{y},\mathcal{A}(\bm{z}))$ starting from $\bm{z}^{\prime}_{i+1}$
@@
5:  for $i=0$ to $T-1$ do $\triangleright$ whole integration path
@@
10:  $\bm{z}^{(e+1)}\leftarrow\bm{z}^{(e)}-\eta^{(e)}\cdot\nabla_{\bm{z}}\mathcal{L}(\bm{z}_{i+1})$triangleright reducing loss
@@ 2.3 Foundation FM priors for IPs
Table 1: A comparison between foundation, domain-specific, and untrained priors for Gaussian deblurring on AFHQ (resolution: $\times 256$ 256; DS: domain-specific; FD: foundation). Bold: best, & underline: second best, for each metric/column
@@
DIP 32.9470 0.9075 0.1873 0.5024
D-Flow (DS) 30.2512 0.8553 0.1183 0.6536
D-Flow (FD) 30.0975 0.8861 0.2270 0.5404
The availability of large-scale training data has enabled the development of foundation generative models for several domains, most of them based on FM (or DM) models, e.g., stable diffusion (v3) [15] and FLUX [14] for images, Sora [22] and Veo for videos, Cosmos world model [16]. So, recent IP methods based on pretrained FM priors have started to shift from domain-specific priors to these foundation priors.
However, while these foundation priors only constrain the object to be physically meaningful (e.g., the object being a natural image), domain-specific priors provide much more semantic and perhaps structural information about the object (e.g., the object being a facial or brain MRI image), i.e., foundation priors alone are considerably weaker than domain-specific priors. In fact, untrained priors such as deep image priors may be powerful enough to promote physically meaningful reconstruction [7, 8, 9, 10, 11]. A quick comparison summarized in Table 1 confirms our intuition: recent IP methods with foundation priors perform worst compared to domain-specific and even untrained priors.
To strengthen the object prior, [19] assumes similarity of $\bm{x}$ and $\bm{y}$—valid for typical image restoration problems, and initializes Eq. 2.3 with
@@
3 Method
We focus on image restoration with foundation FM priors, and aim to strengthen the image prior based on (1) the similarity of $\bm{x}$ and $\bm{y}$, and (2) Gaussianity in FM priors, in the same vein as [19] but improving upon theirs. Adding text prompts as in [17] is presumably helpful, but the reliable generation of high-quality prompts for degraded images seems tricky; we do not pursue it here for leave it for future work.
3.1 Time-dependent warm-up strategy
In the standard FM setting, the source distribution $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, whereas the initialized $\bm{z}_{0}$ in Eq. 2.4 has a distribution $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$. One might not worry about this distribution mismatch, as both are supported on the entire ambient space anyway. But finite-sample training in practice causes the gap: due to the famous concentration of measure phenomenon [23], virtually all training samples drawn from $\mathcal{N}(\bm{0},\bm{I})$ come from an ultra-thin shell around $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$, a sphere centered at $\bm{0}$ and with a radius $\sqrt{d}$: $\mathbb{P}[|\norm{\bm{z}}_{2}-\sqrt{d}|\geq t]\leq 2e^{-ct^{2}}$ for a universal constant $c>0$. If we write that ultra-thin shell spanned by the finite training set as $\mathcal{S}$, the generation function $\mathcal{G}_{\bm{\theta}}$ is effectively trained on inputs from the domain $\mathcal{S}$, not the entire ambient space: the behavior of $\mathcal{G}_{\bm{\theta}}$ on $\mathcal{S}^{c}$ is largely undetermined. Now, samples from $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$ concentrate around another ultra-thin shell around $\mathbb{S}^{d-1}(\sqrt{\alpha}\bm{y}_{0},\sqrt{(1-\alpha)d})$, which has only a negligibly small intersection with $\mathcal{S}$. So, the initialization in Eq. 2.4 lies in $\mathcal{S}^{c}$ with a very high probability. Given that the behavior of $\mathcal{G}_{\bm{\theta}}$ can be wild on $\mathcal{S}^{c}$, this initialization strategy seems problematic.
We take a different route. We know that a typical flow of FM models takes the form:
$\bm{z}_{t}=\alpha_{t}\bm{x}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.1)
Now, when $\bm{x}$ and $\bm{y}$ are close, $\bm{x}=\bm{y}+\bm{\varepsilon}$ for some small $\bm{\varepsilon}$. So, we can write the flow as
$\bm{z}_{t}=\alpha_{t}(\bm{y}+\bm{\varepsilon})+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ (3.2)
for unknown $\bm{\varepsilon}$. When $\alpha_{t}$ is sufficiently small—i.e., we are sufficiently close to $t=0$ in the flow, $\alpha_{t}\bm{\varepsilon}$ can be negligibly small, leading to the approximate flow
$\bm{z}_{t}\approx\alpha_{t}\bm{y}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.3)
In practice, we do not know how small $\alpha_{t}$ should be, so we leave it learnable. This leads to the following formulation
$)).\displaystyle\boxed{\min\nolimits_{\bm{z},t\in[0,1]}\;\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}\pqty{\alpha_{t}\bm{y}+\beta_{t}\bm{z},t}}}.$ (3.4)
@@ 3.1 Time-dependent warm-up strategy
Due to the approximation errors in matching the ideal flow during FM training, as well as when approximating Eq. 3.2 using Eq. 3.3, the distribution of $\bm{z}_{t}$ could be slightly off the ideal distribution. To rectify this, we perform a lightweight variance calibration in our implementation:
@@
where $\mathrm{Var}(Z_{t})$ is the scalar-variable variance across all dimensions and the generation paths of 512512512 samples, and $\mathrm{Var}(\bm{z}_{t})$ scalar-variable variance for $\bm{z}_{t}$ across all dimensions.

### v2

@@ 1 Introduction
Traditional ideas for solving IPs rely on optimization formulations, often motivated under the Maximum A Posteriori (MAP) estimation principle:
@@
Here, minimizing data fitting loss $\ell(\bm{y},\mathcal{A}(\bm{x}))$ encourages $\bm{y}\approx\mathcal{A}(\bm{x})$, and the regularization term $\Omega(\bm{x})$ encodes prior knowledge of ideal solutions to resolve ambiguities and hence mitigate potential ill-posedness. The optimization problems are often solved by gradient-based iterative methods. Advances in deep learning (DL) have revolutionized IP solving. Different DL-based approaches to IPs operate with different data-knowledge tradeoffs. For example, supervised approaches take paired datasets $\left\{\bm{y}_{i},\bm{x}_{i}\right\}_{i=1}^{N}$ and directly learn the inverse mapping $\bm{y}\mapsto\bm{x}$, with or without using $\mathcal{A}$ [Ongie2020DeepLearning, monga2021algorithm, zhang2024wrong]; alternatively, data-driven priors learned from object-only datasets $\left\{\bm{x}_{i}\right\}_{i=1}^{N}$ can be integrated with Eq. 1.1 [oliviero2025generative, daras2024surveydiffusionmodelsinverse, Wang2024DMPlugAP]; strikingly, untrained DL models themselves can serve as plug-in priors for Eq. 1.1, without any extra data [alkhouri2025understanding, wang2023early, li_deep_2023, zhuang_blind_2023, zhuang_practical_2023, li2021self]. [Ongie2020DeepLearning] give comprehensive reviews of these ideas.
In this paper, we focus on solving IPs with pretrained deep generative priors (DGPs) [oliviero2025generative]. Compared to supervised approaches that perform task-specific training, these approaches enjoy great flexibility, as DGPs can be plugged in and reused for different tasks related to the same family of objects. Among the different DGPs, we are most interested in those based on diffusion models (DMs), and the more general flow-matching (FM) models [lipman2024flow], as they are the backbone of state-of-the-art (SOTA) generative models in several domains [labs2025flux1kontextflowmatching, esser2024scalingrectifiedflowtransformers, Agarwal2025Cosmos].
Numerous recent works target solving IPs with pretrained DM/FM priors [daras2024surveydiffusionmodelsinverse]. Although promising, most of them are based on domain-specific DM/FM priors, e.g., trained on the FFHQ dataset for human faces and the LSUN bedrooms dataset for bedroom scenes. This limits the practicality of these methods, as domain-specific DM/FM models are not always readily available. On the other hand, the emergence of domain-agnostic foundation DM/FM models, such as Stable Diffusion [esser2024scalingrectifiedflowtransformers] and Flux [labs2025flux1kontextflowmatching] for images, seems to obsolete domain-specific developments; [kim2025flowdpsflowdrivenposteriorsampling, patel2024steering, ben2024d] propose such ideas. However, the reported performance clearly lags behind those obtained with domain-specific FM/DM priors, and even behind those obtained with untrained priors; see Section 2.3. This should not be surprising, as in terms of solving any particular IP, foundation priors are considerably weaker than domain-specific ones and not clearly stronger than untrained priors.
@@ 2.1 Flow matching (FM)
Flow Matching (FM) models are an emerging class of deep generative models [lipman2024flow]. They learn a continuous flow to transform a prior distribution $p_{0}(\bm{z})$ into a target distribution $p_{1}(\bm{x})$—in the same spirit of continuous normalizing flow (CNF), described by an ordinary differential equation (ODE)
@@
DMs based on probability flow ODEs can also be interpreted as FMs, although they match the score functions induced by the chosen probability path, not the vector field [lipman2024flow]. So, FM can be viewed as a general generative framework that covers DM as well besides other algorithmic possibilities.
@@ 2.2 Pretrained FM priors for IPs
Recent methods that use pretrained FM priors for solving IPs can be classified into two families: (1) The interleaving approach interleaves the generation steps (i.e., numerical integration steps) with gradient steps toward feasibility (i.e., making $\bm{x}$ satisfies that $\bm{y}\approx\mathcal{A}(\bm{x})$[pokle2023training, kim2025flowdpsflowdrivenposteriorsampling, patel2024steering, martin2025pnpflowplugandplayimagerestoration]; see a sample algorithm in Algorithm 1. Despite the simplicity and empirical effectiveness on simple IPs, these methods might not converge or return an $\bm{x}$ that respects the data prior (i.e., manifold feasibility) and/or satisfies the measurement constraint $\bm{y}\approx\mathcal{A}(\bm{x})$ (i.e., measurement feasibility); and (2) the plug-in approach views the entire generation process (lines 4–8 in Algorithm 2) as a function $\mathcal{G}_{\bm{\theta}}$ that maps any source sample to a target sample, and plugs the prior into Eq. 1.1 to obtain a unified optimization formulation [ben2024d, Wang2024DMPlugAP, wang2025temporal] ($\circ$: function composition):
[…]
@@
5:  $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$ $\triangleright$ learned velocity
6:  $\bm{z}^{\prime}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$ $\triangleright$ discrete integration
7:  $\bm{z}_{i+1}\leftarrow$ $(\bm{y},\mathcal{A})$-driven update of $\bm{z}^{\prime}_{i+1}$ $\triangleright$ reducing $\ell(\bm{y},\mathcal{A}(\bm{z}))$ starting from $\bm{z}^{\prime}_{i+1}$
@@
5:  for $i=0$ to T1T-1 do $\triangleright$ whole integration path
@@
10:  $\bm{z}^{(e+1)}\leftarrow\bm{z}^{(e)}-\eta^{(e)}\cdot\nabla_{\bm{z}}\mathcal{L}(\bm{z}_{i+1})$ $\triangleright$ reducing loss
@@ 2.3 Foundation FM priors for IPs
Table 1: A comparison between foundation, domain-specific, and untrained priors for Gaussian deblurring on AFHQ-Cat (resolution: $256\times 256$; DS: domain-specific; FD: foundation). Bold: best, & underline: second best, for each metric/column.
@@
DIP 27.5854 0.7179 0.3898 0.2396
D-Flow (DS) 28.1389 0.7628 0.2783 0.5871
D-Flow (FD) 25.1453 0.6829 0.5213 0.3228
FlowDPS (DS) 22.1191 0.5603 0.3850 0.5417
FlowDPS (FD) 22.1404 0.5930 0.5412 0.2906
The availability of large-scale training sets has recently fueled intensive development of foundation generative models in several domains, most of them based on FM models and variants, e.g., Stable Diffusion V3 (and newer) [esser2024scalingrectifiedflowtransformers] and FLUX.1 [labs2025flux1kontextflowmatching] for images, OpenAI Sora [OpenAISora2024] and Google Veo [Veo_DeepMind_2025] for videos, and Nvidia Cosmos world model [Agarwal2025Cosmos]. By contrast, domain-specific FM models are not always readily available (e.g., due to the lack of training data for scientific applications). So, recent IP methods based on pretrained FM priors have started to explore foundation priors.
However, while these foundation priors only constrain the object to be physically meaningful (e.g., the object being a natural image), domain-specific priors provide much more semantic and perhaps structural information about the object (e.g., the object being a facial or brain MRI image), i.e., foundation priors alone are considerably weaker than domain-specific priors. In fact, untrained priors such as deep image priors may be powerful enough to promote physically meaningful reconstruction [alkhouri2025understanding, wang2023early, li_deep_2023, zhuang_blind_2023, zhuang_practical_2023]. A quick comparison summarized in Table 1 confirms our intuition: recent IP methods with foundation priors perform worst compared to domain-specific and even untrained priors.
To strengthen the object prior, [ben2024d] assumes similarity of $\bm{x}$ and $\bm{y}$—valid for typical image restoration problems, and initializes Eq. 2.3 with
@@
where $\bm{y}_{0}$ is the backward solution of the governing ODE, i.e., $\bm{y}_{0}=\bm{y}+\int_{1}^{0}\bm{v}_{\bm{\theta}}(\bm{y}_{t},t)dt$, or inverted seed in other words. Moreover, they promote the Gaussianity of the seed $\bm{z}_{0}$ by recognizing that $\norm{\bm{z}_{0}}_{2}^{2}$ follows a $\chi^{2}$ distribution and thus regularizes its negative log-likelihood; [kim2025flowdpsflowdrivenposteriorsampling] takes automatically generated text prompts for $\bm{y}$ from another model as text conditions, as all recent foundation FM models allow text-prompted generation.
3 Method
We focus on image restoration with foundation FM priors, and aim to strengthen the image prior based on (1) the similarity of $\bm{x}$ and $\bm{y}$, and (2) Gaussianity in FM priors, in the same vein as [ben2024d] but improving upon theirs.
3.1 Time-dependent warm-up strategy
In the standard FM setting, the source distribution $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, whereas the initialized $\bm{z}_{0}$ in Eq. 2.4 has a distribution $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$. One might not worry about this distribution mismatch, as both are supported on the entire ambient space anyway. But finite-sample training in practice causes the gap: due to the famous concentration of measure phenomenon [vershynin2018high], virtually all training samples drawn from $\mathcal{N}(\bm{0},\bm{I})$ come from an ultra-thin shell around $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$, a sphere centered at $\bm{0}$ and with a radius $\sqrt{d}$: $\mathbb{P}[|\norm{\bm{z}}_{2}-\sqrt{d}|\geq t]\leq 2e^{-ct^{2}}$ for a universal constant $c>0$. If we write that ultra-thin shell spanned by the finite training set as $\mathcal{S}$, the generation function $\mathcal{G}_{\bm{\theta}}$ is effectively trained on inputs from the domain $\mathcal{S}$, not the entire ambient space: the behavior of $\mathcal{G}_{\bm{\theta}}$ on $\mathcal{S}^{c}$ is largely undetermined. Now, samples from $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$ concentrate around another ultra-thin shell around $\mathbb{S}^{d-1}(\sqrt{\alpha}\bm{y}_{0},\sqrt{(1-\alpha)d})$, which has only a negligibly small intersection with $\mathcal{S}$. So, the initialization in Eq. 2.4 lies in $\mathcal{S}^{c}$ with a very high probability. Given that the behavior of $\mathcal{G}_{\bm{\theta}}$ can be wild on $\mathcal{S}^{c}$, this initialization strategy seems problematic.
We take a different route. We know that a typical flow of FM models takes the form
$\bm{z}_{t}=\alpha_{t}\bm{x}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}),$ (3.1)
where $\alpha_{t}$ and $\beta_{t}$ are known functions of $t$ with the property that
$\begin{cases}\alpha_{t}{\scriptstyle\searrow}0,\beta_{t}{\scriptstyle\nearrow}1\;\text{as}\;t\to 0\\ \alpha_{t}{\scriptstyle\nearrow}1,\beta_{t}{\scriptstyle\searrow}0\;\text{as}\;t\to 1,\end{cases}$ (3.2)
where $\scriptstyle\nearrow$ and $\scriptstyle\searrow$ indicate monotonically increasing and decreasing, respectively. Now, when $\bm{x}$ and $\bm{y}$ are close, $\bm{x}=\bm{y}+\bm{\varepsilon}$ for some small (i.e., $\norm{\bm{\varepsilon}}$ is small compared to $\norm{\bm{x}}$ and $\norm{\bm{z}}$) but unknown $\bm{\varepsilon}$. So, we can write the flow as
$\bm{z}_{t}=\alpha_{t}(\bm{y}+\bm{\varepsilon})+\beta_{t}\bm{z}=\alpha_{t}\bm{y}+\beta_{t}\bm{z}+\alpha_{t}\bm{\varepsilon}$
$\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.3)
To eliminate the unknown $\bm{\varepsilon}$, we can approximate the exact flow in Section 3.1 with the following approximate flow
$\bm{z}_{t}\approx\alpha_{t}\bm{y}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ (3.4)
with an approximation error $\alpha_{t}\bm{\varepsilon}$. To control the error, (1) if $\bm{\varepsilon}$ is relatively large, a small $\alpha_{t}$ is desirable; (2) if $\bm{\varepsilon}$ is already relatively small, a relatively large $\alpha_{t}$ is acceptable. So, although we do not know $\bm{\varepsilon}$ itself and hence its magnitude, with appropriate $\alpha_{t}$ we can always make $\alpha_{t}\bm{\varepsilon}$ sufficiently small. So, we leave $\alpha_{t}$ as a learnable parameter. Since $\alpha_{t}$ is a known function of $t$, we simply need to leave $t\in[0,1]$ learnable, leading to our warm-start formulation
@@ 3.1 Time-dependent warm-up strategy
Due to the approximation errors in matching the ideal flow during FM training, as well as when approximating Section 3.1 using Eq. 3.4, the distribution of $\bm{z}_{t}$ could be slightly off the ideal distribution. To rectify this, we perform a lightweight variance calibration in our implementation:
@@
where $\mathrm{Var}(Z_{t})$ is the scalar-variable variance across all dimensions and the generation paths of $512$ samples, and $\mathrm{Var}(\bm{z}_{t})$ scalar-variable variance for $\bm{z}_{t}$ across all dimensions.